# Instruct, Not Assist: LLM-based Multi-Turn Planning and Hierarchical Questioning for Socratic Code Debugging

### v3

@@ 1 Introduction
With the rapidly expanding conversational and reasoning abilities of large language models (LLMs), there has been a substantial rise in demand for exploiting their capabilities within a multitude of educational applications Kasneci et al. (2023) in order to widen accessibility via personalized feedback. Specifically, several recent works explore the use of LLMs for providing feedback and guidance to students Wang et al. (2023); Kazemitabaar et al. (2024); Sheese et al. (2024); Lyu et al. (2024). However, LLMs are typically optimized to generate customer-serving, assistant-like responses, which also translates into the types of questions asked. Especially for educational domains, this style of questioning can be sub-optimal Cotton (1988); Sahamid (2016); Yang et al. (2005); Wilson (1987). For instance, if a student is seeking help from an instructor for correcting their mistakes (e.g., debugging their buggy code), we consider two forms of potential responses: assistant-like and instructor-like. As shown in Figure 1, an assistant-like response would not be a successful educational interaction as it leads to the Assistant directly providing an answer. On the other hand, an Instructor-like response reflects the educational philosophy of Socratic questioning.
@@
    Tree-based Socratic questioning: An Instructor generates turn-level Socratic questions conditioned on both the Student’s current knowledge state and misunderstanding(s), the latter derived from their responses to the Instructor’s questions. This step dynamically constructs a Socractic question tree.
@@
  • We construct a novel challenging multi-bug debugging dataset with 150 expert-annotated conceptual and syntactical problems and buggy solutions/fixes.
  • Extensive experiments on an existing benchmark and our constructed dataset demonstrate that TreeInstruct can be universally applied to both open and closed source-settings.
  • We also showcase that TreeInstruct’s strong Socratic questioning abilities widely outperform all baselines through both (1) rigorous quantitative and qualitative expert evaluation (on average, preferred over 78.43% of the time) and (2) real-world interactions with students of varying coding abilities.
@@ 2.3 LLMs for Interactive Education
Recent generative approaches within the AI tutoring space have attempted to generate responses which cater to the student’s type of mistake or request, but only in single-turn settings. CodeAid Kazemitabaar et al. (2024) is an assistive tool that helps students debug their code. In their "Help Fix Code" and "Question from Code" modules, the Instructor provides single-turn responses to the Student for answering questions, explaining concepts, and helping to write code. In contrast, TreeInstruct aims to instruct the Student socratically through questions. It also takes into account the inherent multi-turn nature of conversational conversations during the "planning" stage. BRIDGE Wang et al. (2023) is an Instructor-like framework that aims to help students with math mistakes. The LM estimates the type of error, the strategy of error remediation, and the instructor intention behind the remediation (all are chosen from a predetermined set). However, our methodology makes use of a more structured approach that works well for multi-turn conversations while adapting to the Student’s learning.
@@
Figure 3: We detail the process for tree-based question generation. Blue and orange text/backgrounds indicate that the Instructor and Verifier are performing the task respectively.
@@ 3 Methodology
As shown in Figure 2, TreeInstruct aims to dynamically guide the multi-turn conversation based on its estimated state space. Section 3.1.2 provides an overview of the three different agents we use and their respective roles during the state space generation/update and tree construction processes (outlined in Figure 3). This allows TreeInstruct to respond to the Student’s current level of understanding adequately. Algorithm 1 contains the pseudo-code for all components in our method.
@@ 3.1.1 Problem Description
As input, the Instructor is given the Student’s buggy code that contains $e$ errors, a problem statement, bug descriptions, and their respective fixes. The Instructor guides the Student to generate a list of all bug fixes based on their interactions with the Instructor. The overall goal is for the Student to resolve their own conceptual and syntactical errors in a Socratic fashion to reach the correct code.
@@ 3.1.2 Agents
In a real-world setting, a Socratic educator (e.g., an instructor, a teaching assistant, a professor) executes two tasks when interacting with a Student: (1) ask relevant questions to the Student, and (2) assess the Student’s understanding based on their responses. Following this cyclical pattern, we break down our educator into two roles: an Instructor and a Verifier, with persona prompts specified in Tables 9 and 11 in Appendix I, respectively. The Instructor and Verifier perform their respective tasks specified in Algorithm 1 via zero and one-shot prompting. The Instructor agent’s job is to generate questions to ask the Student (GenerateQuestion, GenerateSiblingQuestion, and GenerateChildQuestion in Alg. 1; details provided in Section 3.3). The Verifier agent has a significantly more involved role:
@@
    State space estimation (Section 3.2): The Verifier determines a set of tasks which will lead a Student’s to understanding and correcting their problem and buggy code. This is GenerateState in Alg. 1.
@@ 3.2 State Space Estimation
We define the state space as the set of all possible tasks that a Student could perform to correct their buggy code. We claim that the optimal state space can be represented by a series $S$ of $k$ tasks which leads the Student from their buggy code $B$ to (1) understanding their conceptual and syntactical mistakes and (2) correcting their code. Each of these tasks is a state variable $\tau_{i}$ which either has a value of True or False based on whether or not the Student has completed it. At the very beginning of the Instructor-Student conversation, all of these variables are set to “False”. We provide the estimated state space used in Figures 2 and 3.
@@
The state variables $\tau_{i}$ are structured such that earlier tasks have a higher priority, as their completion may consequently resolve later tasks. For instance, a student’s buggy code may reflect that they do not understand an edge case mentioned in the problem statement. However, once this misunderstanding is resolved, the Student may simultaneously correct their related syntactical mistakes. On the other hand, attempting to resolve the syntactical mistakes, “Modifying the condition in the if statement”, beforehand may lead to an unproductive and less structured conversation overall.
@@ 3.3 Tree-Based Questioning
The Instructor conditionally generates sibling questions at level $l$ if and only if the Student incorrectly answers the Instructor question (lines 6 and 10 in Alg. 1). As shown in the second and third question of Figure 3, these questions must lead to the same level of target understanding as the original generated question intended so therefore, the question can be rephrased or made more specific. To ensure this, we ground the question generation based on two things: (1) the previous questions from level $l$, and (2) the Verifier’s explanation for why the Student got the question wrong.
@@ 3.4 Adaptive Conversation Restructuring
Once the Verifier agent determines that the target state/task has been resolved, we exploit the same process to update all remaining tasks $\tau\in S$, as multiple dependent bugs may have been concurrently resolved within the same tree. After at least the target state variable has been resolved (line 13 in Alg. 1), we create a new tree for any remaining tasks, as shown in the first interaction of Figure 3. This step is crucial to the multi-bug setting, as mutually independent bugs would benefit from having separate and distinct trees of questioning.
@@
  • Bug fixes: After a task $\tau$ has been resolved, the Student is prompted to provide a list of natural language bug fixes (e.g. "Replace ‘i‘ with ‘i+1‘ on line 6.") based on their entire conversation history with the Instructor. The Verifier will determine if all of the ground-truth bug fixes have an isomorphic counterpart within the set of suggested Student bug fixes. Isomorphism can be defined as (1) having the same conclusion or output, (2) sharing the same underlying logical structure or pattern, and/or (3) being convertible to each other through a series of logical transformations. If all ground-truth bug fixes have been resolved, then we may stop early.
@@
Table 1: Results on the Socratic Debugging Benchmark Dataset (Single Bug). Bolded and $\dagger$ values denote the top 2 methods respectively.
@@
Methods Avg. Turns Success Relevant Indirect Logic Success Relevant Indirect Logic
@@ 4.1 Experimental Setup
In order to evaluate TreeInstruct, we utilize a proxy Student agent based on the Mistral-7B-Instruct model Jiang et al. (2023) to mimic the abilities of a student while responding to the Instructor. The prompt we use to define the Student persona is outlined in Table 10 of Appendix I. We additionally provide GPT4 API experimental set up details in Appendix G.
Table 2: Results on the MULTI-DEBUG dataset. In total, 1-bug has 29 syntactical and 21 conceptual bugs, 2-bug has 50 syntactical and 50 conceptual bugs, and 3-bug has 78 syntactical and 72 conceptual bugs. Bolded and $\dagger$ values denote the top 2 methods respectively.
@@
Bugs Methods Avg. Turns Success Relevant Indirect Logic Success Relevant Indirect Logic
@@
4.2 Datasets
We use two datasets to evaluate our method on. First, the Socratic Debugging Benchmark dataset from Al-Hossami et al. (2023b), which consists of 149 problems, each with a problem statement, student buggy code, bug fixes and descriptions in English, and correct code. Each problem has one syntactical or conceptual bug. Second, to challenge our method, we also craft a novel dataset, MULTIDEBUG, based on 50 popular programming problems [^2] https://github.com/Garvit244/Leetcode. For each of the 50 problems, we inject 1, 2, and 3 bug(s) that a student would make for a total of 150 different samples. We keep track of these bugs with matching bug fixes and descriptions.
Bugs are either conceptual or syntactical. Conceptual bugs usually cause runtime errors or result in incorrect output. Examples include misunderstanding the problem statement, encountering an infinite loop, or incorrectly using a library or mathematical operator (/// vs ///// / in Python). Syntactical bugs cause compilation errors due to incorrect Python syntax (e.g., missing a colon).
4.3 Baselines
To determine the success of TreeInstruct, we also measure the performance on a few baselines. First is a baseline called Vanilla. Given the same input as TreeInstruct’s Instructor, this method simply asks the base model to ask Socratic questions to the Student - it does not utilize the tree structure, nor does it estimate the Student’s knowledge. We use both Meta-Llama-3-8B-Instruct [Uncaptioned image]Touvron et al. (2023) and GPT-4 [Uncaptioned image]Achiam et al. (2023) as base models for the Vanilla baseline.
Second, we use BRIDGE Wang et al. (2023). Since we are adapting this for Socratic code debugging, we use the error type, the remediation strategy, and the remediation intention to guide the question generation, along with the problem-specific input given to TreeInstruct’s Instructor. For both baselines, we limit the conversations to 20 turns per number of bugs.
@@ 4.4 Evaluation Metrics:
We perform qualitative and quantitative evaluation of our methods. Details for each metric are provided in Appendix A. The scores are averaged across all turns and then averaged across all problems. In the results, we scale the scores by 100.
Qualitative Metrics:
We develop a binary scale to assess the Socratic quality of questions. We measure each metric manually, giving a score of 1 if the attribute is met, and 0 otherwise.
Relevance (Relevant) measures whether the instructor’s question was pertinent to the errors in the Student’s code. Indirectness (Indirect) measures if the instructor’s question refrained from directly revealing solutions to the bugs. Finally, Logical Flow (Logic) checks if the instructor’s question promoted a coherent conversation, facilitating the Student’s problem-solving process. Each Instructor question is assigned with a binary value for each of the three attributes.
Quantitative Metrics:
We apply quantitative metrics to objectively evaluate the effectiveness and efficiency of our framework. We calculate the overall Success Rate (Success) with the number of bug fixes generated by the Student that are isomorphic to the ground truth set of bug fixes. We also compute the average number of turns (Avg. Turns) required by the method to reach the goal state.
4.5 Overall Results
In Tables 1 and 2, we see that with more structured representations of student knowledge and conversation state, TreeInstruct demonstrates significant improvements beyond the baselines. Across all multi-bug settings, we see an overall improvement of 16.6% and 11.59% in the success rates for syntactical and conceptual bugs, respectively. We also see an improvement of 13.47% and 14.89% for syntactical and conceptual bugs, respectively, across the three conversation metrics. In the 1-bug setting, we see that the Vanilla [Uncaptioned image]baseline has the highest success for conceptual bugs. However, this setting simultaneously has the lowest Indirectness score, indicating that questions were very direct, and gave hints towards the bug fixes, which evidently increased the success rate. We see the same trend in the syntactical, single bug on Vanilla [Uncaptioned image]setting in Table 1. Overall, TreeInstruct demonstrates strong performance despite drastically different base models, [Uncaptioned image]and [Uncaptioned image].
Side-by-Side evaluation:
Using the conversational metrics, we performed a side-by-side evaluation that measures how often a user prefers our method TreeInstruct over the baselines. More details are in Appendix E, in which we see that on average, TreeInstruct was preferred over BRIDGE 79.43% more, and over Vanilla 77.43% more.
Human Student Interaction:
We also conduct a separate case study where human students directly interact with TreeInstruct (details in Appendix F). We see that with varying levels of programming backgrounds, TreeInstruct is able to help all students resolve the bugs in their code.
4.5.1 Analysis
Conceptual bugs are easier to solve than syntactical.
In both tables, across all settings, it can be seen that questions targeted towards conceptual bugs have higher scores than those towards syntactical bugs. Syntactical bugs might be "harder to see" for the language model as it goes against the generation process to generate syntactically incorrect code. Breaking it down, a language model trained to generate code will always add a colon at the end of for loops, if-statements, and method signatures because the model is trained to do that. Even though the buggy code might have a missing colon, the language model might ignore it.
More bug-specific state variables helps (1) generate more relevant questions and (2) maintain conversational flow.
Table 2 shows that scores for Relevance and Logic on conceptual bugs decreases as the number of bugs increases (from 100% relevance and 94.76% logical flow in 1-bug to 98.4% relevance and 93.63% logical flow in 3-bug). The state space representation for the 1-bug setting is much less compact as that of the 2- and 3-bug settings. To elaborate, 1-bug state space representations suggest 3 or 4 state variables (more subtasks) to solve a bug that 2- or 3-bug state space representations take 1 or 2 state variables to solve (example provided in Appendix D). This indicates that, across states, 1-bug settings have an easier time keeping questions relevant to the bugs (as there is only one). Additionally, the conversation flows better as there are no inter-bug dependencies that the Instructor has to take into account.
Bug dependency affects success rate.
Table 2 shows that the success rate (SR) for 3-bug and 1-bug are higher than that of 2-bug. 1-bug settings are overall relatively easier given that only one bug must be resolved. However, compared to the baselines, which feature low success and logic rates, Tree-Instruct demonstrates a comparatively strong performance in 3-bug settings. This is likely due to its state space and tree creation structure factoring in the inter-bug dependencies.
For example, in the Fibonacci problem in Figure 1, a student could have made the following two bugs: (1) they did not add a base case for the recursion, and (2) they did not correctly write the recursive call. Once they solve one of the bugs, they will have understood recursion better, enabling them to solve the other bug easily. TreeInstruct’s prioritization of conceptual errors in the state space estimation (Section 3.2) and dependency awareness (Section 3.4) are the key to its high 3-bug performance. The same cannot necessarily be said for the 2-bug setting as it could have mutually independent bugs that require special attention to solve.
5 Conclusion
This paper proposes a novel method, TreeInstruct, for state space estimation and dynamic tree-based questioning for multi-turn Socratic instruction. We construct a novel multi-bug debugging dataset, MULTI-DEBUG, with 150 expert-annotated conceptual and syntactical problems and buggy solutions/fixes. Extensive experiments on an existing benchmark and MULTI-DEBUG demonstrate that TreeInstruct can be universally applied to both open and closed source-settings. We also showcase that TreeInstruct’s strong Socratic questioning abilities widely outperform all baselines through both (1) rigorous quantitative and qualitative expert evaluation (preferred over 77.94% of the time), and (2) real-world interactions with students of varying coding abilities (in Appendix F).
6 Limitations & Future Work
Moreover, the base language models cannot fully grasp abstract concepts such as trees and linked lists. Even humans require diagrams to work out potential solutions or teach one another. With the language-reliant teaching strategies, our method might not be able to effectively teach in these domains.
These limitations give way to exciting future work. Firstly, we can make use of vision language models to provide students with multi-modal teaching strategies, instead of relying solely on language. Additionally, we can enhance the framework, so it will explore new instruction methods when the questioning becomes cyclical. This can also help make the Instructor more reliable to generate consistent output across multiple runs on the same problem. Furthermore, we can utilize a structured fine-tuning approach to help the model better leverage the Verifier feedback and tree-based question generation process and make hierarchical Socratic planning and questioning inherent to a model. Overall, TreeInstruct can also be extended to automatically generalize to different teaching domains (e.g., quantitative reasoning).
7 Ethics Statement
8 Acknowledgements
Appendix A Evaluation Metrics
Here, we describe our qualitative and quantitative metrics in depth. The scores are averaged across all turns and then averaged across all problems of the same setting and dataset. In our results, we scale the scores by 100.
@@ Appendix A Evaluation Metrics
We develop a binary scale to assess the Socratic quality of questions. Previous work identifies multiple dimensions of Socratic questioning, including relevance to specific needs, implicitness of the answer, and structural coherence. For each question, we measure the below attributes of the conversation manually (giving a score of 1 if the attribute is met, and 0 otherwise):

### v4

@@ 1 Introduction
With the rapidly expanding conversational and reasoning abilities of large language models (LLMs), there has been a substantial rise in demand for exploiting their capabilities within a multitude of educational applications Kasneci et al. (2023) in order to widen accessibility via personalized feedback. Specifically, several recent works explore the use of LLMs for providing feedback and guidance to students Wang et al. (2023); Kazemitabaar et al. (2024); Sheese et al. (2024); Lyu et al. (2024). However, LLMs are typically optimized to generate customer-serving, assistant-like responses, which also translates into the types of questions asked. Especially for educational domains, this style of questioning can be suboptimal Cotton (1988); Sahamid (2016); Yang et al. (2005); Wilson (1987). For instance, if a student is seeking help from an instructor for correcting their mistakes (e.g., debugging their buggy code), we consider two forms of potential responses: assistant-like and instructor-like. As shown in Figure 1, an assistant-like response would not be a successful educational interaction, as it leads to the Assistant directly providing an answer. On the other hand, an Instructor-like response reflects the educational philosophy of Socratic questioning.
@@
    Tree-based Socratic questioning: An Instructor generates turn-level Socratic questions conditioned on both the Student’s current knowledge state and misunderstanding(s), the latter derived from their responses to the Instructor’s questions. This step dynamically constructs a Socratic question tree.
@@
  • We construct a novel multi-bug debugging dataset with 150 expert-annotated, challenging conceptual and syntactical bugs and their fixes.
  • Extensive experiments on an existing benchmark and our constructed dataset demonstrate that TreeInstruct can be universally applied to both open and closed source-settings. We also showcase that TreeInstruct’s strong Socratic questioning abilities widely outperform all baselines through both (1) rigorous quantitative and qualitative expert evaluation (on average, preferred 78.43% of the time; Student fixes code 24.55% more) and (2) real-world interactions with students of varying coding abilities.
@@ 2.3 LLMs for Interactive Education
Recent generative approaches within the AI tutoring space have attempted to generate responses which cater to the student’s type of mistake or request, but only in single-turn settings. CodeAid Kazemitabaar et al. (2024) is an assistive tool that helps students debug their code. In their "Help Fix Code" and "Question from Code" modules, the Instructor provides single-turn responses to the Student for answering questions, explaining concepts, and helping to write code. However, these modules direct the Student towards where their mistake is and uses natural language to describe the bug fixes. In contrast, TreeInstruct aims to instruct the Student socratically through questions, such that even in natural language, the bug fixes are not provided. BRIDGE Wang et al. (2023) is an Instructor-like framework that aims to help students with math mistakes. The LM estimates the type of error, the strategy of error remediation, and the instructor intention behind the remediation (all are chosen from a predetermined set). However, our methodology makes use of a more structured planning approach that accounts for the inherent multi-turn nature of educational guidance.
@@
Figure 3: We detail the process for tree-based question generation. Blue and orange text/backgrounds indicate that the Instructor and Verifier are performing the task, respectively.
@@ 3 Methodology
As shown in Figure 2, TreeInstruct aims to dynamically guide the multi-turn conversation based on its estimated state space. Section 3.1.2 provides an overview of the three different agents we use and their respective roles during the state space generation/update and tree construction processes (outlined in Figure 3). This allows TreeInstruct to respond to the Student’s current level of understanding adequately. Algorithm 1 contains the pseudocode for all components in our method.
@@ 3.1.1 Problem Description
As input, the Instructor is given the Student’s buggy code that contains $e$ errors, a problem statement, bug descriptions, and their respective fixes. The Instructor guides the Student to generate a list of all bug fixes based on their interactions with the Instructor. The overall goal is for the Student to resolve their own conceptual and syntactical errors in a Socratic fashion to reach the correct code. Note that we assume bug fixes are provided, a common scenario in educational settings (e.g., assignments, exams) where ground-truth solutions are available from human instructors or platforms like LeetCode. As our focus is steering an LLM towards Socratic guidance, generating these solutions for real-world tutoring applications is left for future work.
@@ 3.1.2 Agents
In a real-world setting, a Socratic educator (e.g., an instructor, a teaching assistant) executes two tasks when interacting with a Student: (1) ask relevant questions to the Student, and (2) assess the Student’s understanding based on their responses. Following this cyclical pattern, we break down our educator into two roles: an Instructor and a Verifier, with persona prompts specified in Tables 12 and 14 in Appendix I, respectively. The Instructor and Verifier perform their respective tasks specified in Algorithm 1 via zero and one-shot prompting. The Instructor agent’s job is to generate questions to ask the Student (GenerateQuestion, GenerateSiblingQuestion, and GenerateChildQuestion in Alg. 1; details provided in Section 3.3). The Verifier agent has a significantly more involved role:
@@
    State Space Estimation (Section 3.2): The Verifier determines a set of tasks which will lead a Student to understanding and correcting their problem and buggy code. This is GenerateState in Alg. 1.
@@ 3.2 State Space Estimation
We define the state space as the set of all possible tasks that a Student could perform to correct their buggy code. We claim that the optimal state space can be represented by a series $S$ of $k$ tasks which leads the Student from their buggy code $B$ to (1) understanding their conceptual and syntactical mistakes and (2) correcting their code. Each of these tasks is a state variable $\tau_{i}$ which either has a value of True or False based on whether the Student has completed it. At the very beginning of the Instructor-Student conversation, all of these variables are set to “False”. We provide the estimated state space used in Figures 2 and 3.
@@
The state variables $\tau_{i}$ are structured such that earlier tasks have a higher priority, as their completion may consequently resolve later tasks. For instance, a student’s buggy code may reflect that they do not conceptually understand the definition of the Fibonacci sequence. However, once this misunderstanding is resolved, the Student may simultaneously correct their related syntactical mistakes. On the other hand, attempting to resolve their syntactical mistakes, “Modify the recursive call”, beforehand may lead to an unproductive and less structured conversation overall.
@@ 3.3 Tree-Based Questioning
The Instructor conditionally generates sibling questions at level $l$ if and only if the Student incorrectly answers the Instructor question (lines 6 and 10 in Alg. 1). As shown in the second and third question of Figure 3, these questions must lead to the same level of target understanding as the original generated question intended therefore, the question can be rephrased or made more specific. To ensure this, we ground the question generation based on two things: (1) the previous questions from level $l$, and (2) the Verifier’s explanation for why the Student got the question wrong.
@@ 3.4 Adaptive Conversation Restructuring
Once the Verifier agent determines that the target state/task has been resolved, we exploit the same process to update all remaining tasks $\tau\in S$, as multiple dependent bugs may have been concurrently resolved within the same tree. After at least the target state variable has been resolved (line 13 in Alg. 1), we create a new tree for any remaining tasks, as shown in the first interaction of Figure 3. This step is crucial to the multi-bug setting, as independent bugs would benefit from having separate, distinct trees of questioning.
@@
  • Bug fixes: After a task $\tau$ has been resolved, the Student is prompted to provide a list of natural language bug fixes (e.g. “Replace i with i+1 on line 6.”) based on their entire conversation history with the Instructor. The Verifier will determine if all the ground-truth bug fixes have an isomorphic counterpart within the set of suggested Student bug fixes. Isomorphism can be defined as (1) having the same conclusion or output, (2) sharing the same underlying logical structure or pattern, and/or (3) being convertible to each other through a series of logical transformations [^2] For example, the bug fixes ”Replace if i <= 0 with if i < 0” and ”Check if i is strictly less than 0” are isomorphic because they have the same conclusion and share the same underlying pattern of the mathematical operator $<$.. If all ground-truth bug fixes have been resolved, then we may stop early.
@@
Table 1: Results on Socratic Debugging Benchmark (Single Bug). Bolded and $\dagger$ values denote the top 2 methods.
@@
Methods #T Success Relevant Indirect Logic Success Relevant Indirect Logic
@@ 4.1 Experimental Setup
In order to evaluate TreeInstruct, we utilize a proxy Student agent based on the Mistral-7B-Instruct model Jiang et al. (2023) to mimic the abilities of a student while responding to the Instructor. The prompt we use to define the Student persona is outlined in Table 13 of Appendix I. We additionally provide GPT-4 API experimental setup details in Appendix G.
4.2 Datasets
We evaluate our method on two datasets. First, we use the Socratic Debugging Benchmark dataset from Al-Hossami et al. (2023b), which consists of 149 problems– each with a problem statement, student buggy code, bug fixes and descriptions in English, and correct code. However, these problems lack sufficient difficulty, often requiring small fixes and minimal problem comprehension. To evaluate TreeInstruct on more challenging problems, we craft a novel dataset, MULTI-DEBUG, based on 50 popular programming problems [^3] https://github.com/Garvit244/LeetCode (16 easy, 29 medium, and 5 hard according to their corresponding LeetCode labels). It features longer problems and more involved concepts, requiring TreeInstruct to have more extensive reasoning capabilities for guiding the Student.
For each problem, expert annotators (Appendix D) injected 1, 2, and 3 syntactical or conceptual bug(s) that a typical student would make (a total of 150 different samples for MULTI-DEBUG). Conceptual bugs usually cause runtime errors or result in incorrect output. Examples include misunderstanding the problem statement, encountering an infinite loop, or incorrectly using a library or mathematical operator (/// vs ///// / in Python). Syntactical bugs cause compilation errors due to incorrect Python syntax (e.g., missing a colon). For each bug, we keep track of its fix and description.
4.3 Baselines
We compare TreeInstruct to several baselines and their variants. Given that LLM-based Socratic instruction is a new evolving area with few existing work and no multi-turn methods, we adapt a prompting and an existing single-turn method to our task. The first baseline, Vanilla, is given the same input as TreeInstruct’s Instructor. The base model is prompted to ask Socratic questions to the Student– it does not use any explicit conversational structure or estimate the Student’s knowledge. We use both Meta-Llama-3-8B-Instruct [Uncaptioned image]Touvron et al. (2023) and GPT-4 [Uncaptioned image]Achiam et al. (2023) as base models for the Vanilla baseline.
Table 2: Results on the MULTI-DEBUG dataset. In total, 1-bug has 29 syntactical and 21 conceptual bugs, 2-bug has 50 syntactical and 50 conceptual bugs, and 3-bug has 78 syntactical and 72 conceptual bugs. Bolded and $\dagger$ values denote the top 2 methods, respectively. The bottom two rows are ablation studies performed on 3-Bug setting.
@@ 4.3 Baselines
Bugs Methods #T Success Relevant Indirect Logic Success Relevant Indirect Logic
@@
[Uncaptioned image]No State 16.34 25.51 51.61 97.21 41.09 53.00 47.57 94.70 20.36
[Uncaptioned image]No Teaching 9.69 30.61 90.75 97.61 72.84 50.00 94.62 95.17 68.78
Second, we use BRIDGE Wang et al. (2023). Since we are adapting this for Socratic code debugging, we use their predetermined error types, remediation strategies, and remediation intentions to guide the question generation, along with the problem-specific input given to TreeInstruct’s Instructor. For both baselines, we limit the conversations to 20 turns per number of bugs.
@@ 4.4 Evaluation Metrics:
We perform qualitative and quantitative evaluation of our methods. The scores are averaged across all turns and then averaged across all problems. In the result tables, we scale the scores by 100.
@@
We develop a binary scale to assess the Socratic quality of questions. Al-Hossami et al. (2023b) identifies multiple dimensions of Socratic questioning, including relevance to specific bugs, implicitness of the answer, and structural coherence. For each question, we measure the below attributes of the conversation manually (giving a score of 1 if the attribute is met, and 0 otherwise):